OpenStudy (anonymous):

If f(x) = 2x^2 + 3x and g(x) = x – 2, what is (f + g)(2)? A) 16 B) 14 C) 12 D) 10 E) 8

OpenStudy (noseboy908):

In order to do this problem, you have to add f(x) and g(x) together. Can you restate for me what each of those is?

OpenStudy (anonymous):

how??

OpenStudy (noseboy908):

For the moment, simply restate what you've already given me. What is F(x) defined as, and what is G(x) defined as?

OpenStudy (anonymous):

\[\left( f+g \right)\left( 2 \right)=f \left( 2 \right)+g \left( 2 \right)=?\]

OpenStudy (anonymous):

i do not understand @noseboy908

OpenStudy (anonymous):

i think is B??

OpenStudy (anonymous):

correct.